OpenStudy (whalenjacob):

A pair of equations is shown below: y = 7x − 9 y = 3x − 1 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations.

OpenStudy (whalenjacob):

someone help

OpenStudy (whalenjacob):

@Evoker dude im so sorry my battery died and i couldnt find the chrger

OpenStudy (evoker):

No problem, graphically you would just graph the two equations and see where they cross

OpenStudy (evoker):

The slope is the number next to the x in front of it and the y-intercept is the number after the x.

OpenStudy (whalenjacob):

well they match to 5,2

OpenStudy (whalenjacob):

but how do ianswer the question?

OpenStudy (evoker):

Well for part a they are asking how you would graph it

OpenStudy (evoker):

Haven't seen the other parts but your intersection tell you the solution is when x=5 and y=2

OpenStudy (evoker):

actually you probably had the reverse where x=2 and y=5

OpenStudy (whalenjacob):

ok so would the answer be x=2 and y=5?

OpenStudy (evoker):

for later parts probably, you have only put up part A.

OpenStudy (whalenjacob):

so what would the answer be for this?

OpenStudy (evoker):

This part you would say you draw the two graphs using the methods you did and then find the intersection.

OpenStudy (pythagoras123):

For the line y = mx+c, m is the slope c is the y-intercept set mx+c = 0 to solve for the x-intercept.

OpenStudy (mathmale):

Summary: 1) Graph: y = 7x − 9 y = 3x − 1 2) Determine the coordinates of the intersection of these two lines. Basically, that's it. But you are also asked to identify the slope and y-intercept of each of the given linear functions y = 7x − 9 y = 3x − 1 Hint: For the first function, the slope, m, is 7, and the y-intercept, b, is - 9. What are the slope and y-intercept of the second function?

OpenStudy (mathmale):

You must now describe (in your own words) how to find the solution of this system of linear equations. That "solution" consists of the coordinates of the point at which the two lines intersect.

OpenStudy (mathmale):

This question is tougher than it would be if you were merely asked to find the solution of this system of linear equations. Not only must you know how to find that solution, but also you must be able to explain, in words, in your own words, how to do it. I have given you numerous examples explaining parts of this process.